### v1

# Kernel-U-Net: Hierarchical and Symmetrical Framework for Multivariate Time Series Forecasting

Abstract
Time series forecasting task predicts future trends based on historical information. Recent U-Net-based methods have demonstrated superior performance in predicting real-world datasets. However, the performance of these models is lower than patch-based models or linear models. In this work, we propose a symmetric and hierarchical framework, Kernel-U-Net, which cuts the input sequence into slices at each layer of the network and then computes them using kernels. Furthermore, it generalizes the concept of convolutional kernels in classic U-Net to accept custom kernels that follow the same design pattern. Compared to the existing linear or transformer-based solution, our model contains 3 advantages: 1) A small number of parameters: the parameters size is $O(log(L)^{2})$ where $L$ is the look-back window size, 2) Flexibility: its kernels can be customized and fitted to the datasets, 3) Computation efficiency: the computation complexity of transformer modules is reduced to $O(log(L)^{2})$ if they are placed close to the latent vector. Kernel-U-Net accuracy was greater than or equal to the state-of-the-art model on six (out of seven) real-world datasets.
1 Introduction
Time series forecasting predicts future trends based on recent historical information. It allows experts to track the incoming situation and react timely in critical cases. Its applications range from different domains such as predicting the road occupancy rates from different sensors in the city Lai et al. (2018), monitoring influenza-like illness weekly patient cases Wu et al. (2021), monitoring electricity transformer temperature in the electric power long-term deployment Zhou et al. (2021) or forecasting temperature, pressure and humidity in weather station Liu et al. (2022b) etc.
Over the past few decades, time series forecasting solutions have evolved from traditional statistical methods (e.g., ARIMA Khandelwal et al. (2015) and machine learning techniques (e.g., GBRT Persson et al. (2017)) to deep learning-based solutions, such as RNN recurrent neural networks Tokgöz and Ünal (2018), LSTM Kong et al. (2017), Temporal Convolutional Network (TCN) Hewage et al. (2020) and Transformer-based model Li et al. (2019).
Large-scale time series forecasting addresses the problem of predicting a longer future horizon with a larger look-back window. With the increasing size of data, the cost of learning deep models increases crucially. Moreover, models such as LSTM or RNN suffer gradient vanishing problems in learning long sequences. To balance the quality of prediction and efficiency of learning deep models, researchers proposed hierarchical structure in Duan et al. (2022), pyramidal structure in Pyraformer Liu et al. (2022a), AutoEncoders Asres et al. (2022), U-Net-like model in Yformer Madhusudhanan et al. (2023) or patches in PatchTST Nie et al. (2023). Meanwhile, there is an increasing trend in combining different components for learning particular time series. For example in time series classification, researchers studied hybrid models such as combing GRU Elsayed et al. (2019), LSTM Lin et al. (2020); Khan et al. (2021) or Transformer Huang et al. (2020b) with linear or CNN layers Hewage et al. (2020).
Among the Transformer models applying to time series data, Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and FEDformer Zhou et al. (2022) are the best variants that improve the quality of prediction. Meanwhile, a recent paper Zeng et al. (2023) gives a simple linear model NLinear which outperforms pure transformer-based models on a variety of common benchmarks and challenges the Transformer models for time series forecasting. Later, the authors in PatchTST Nie et al. (2023) argued that applying transformer modules on a point-wise sequence degrades the performance. By adding a linear patch layer, they successfully relieved the overfitting problem of transformer modules and they reached state-of-the-art result on several large datasets.
We remark that models are particularly sensible to different types of datasets. For example, NLinear is more robust against overfitting in small-size datasets such as Etth and EttmZeng et al. (2023) and PatchTST is more expressive in large-size datasets such as Weather and Traffic etc.Nie et al. (2023). At the same time, it is also important to notice the parameter size in Linear models such as NLinear is in $O(L^{2})$ and the computation complexity is $O(L)$. Meanwhile, the parameter size in a transformer block is $O(1)$ but the computational complexity is $O(L^{2})$ in general Transformer modules. It is reduced to $O(Llog(L))$ using Propsparse and to $O(\frac{L}{S}log(\frac{L}{S}))$ using patches, where $S$ is the size of the patch. Based on previous analysis, there is a motivation to propose a more flexible framework for combining the advantages of different modules in high-paced development and testing.
As a classical and highly expressive model in medical image segmentation, Convolutional U-net has an elegant symmetrical encoder and decoder structureRonneberger et al. (2015). It coincides well with the time series forecasting problem because the inputs and outputs are drawn from the same distribution. The first U-Net-like model for time series forecasting was YformerMadhusudhanan et al. (2023), which used Transformers in the encoder and decoder. As mentioned before, applying a Transformer on point-wise data can lead to overfitting. Therefore, we are interested in answering the question of whether there is a U-net-like structure that performs well for the time series forecasting task.
Figure 1: Structure of Kernel U-Net
We propose a general hierarchical and symmetrical framework, Kernel-U-Net, inspired by convolutional U-net Ronneberger et al. (2015) and Yformer Madhusudhanan et al. (2023) for time series forecasting. Kernel-U-Net offers $O(log(L)^{2})$ parameters size for processing a time series of size $L$. Meanwhile, it provides convenience for composing particular models with custom kernels that follow the same design pattern. Since the hierarchical structure reduces the length of the sequence at each level exponentially, the complexity of learning such sequences is simultaneously reduced. In particular, the computation complexity of Transformer and Long Short-term Memory(LSTM) modules is reduced to $O(log(L)^{2})$ and $O(log(L))$ if they are only employed in the layers close to the latent vector. Therefore, Kernel-U-Net helps reduce the computation cost of the Transformer and relieves the gradient vanishing problem in LSTM in learning long sequences.
To fully study the performance and efficiency of Kernel-U-Net, we conduct experiments for time series forecasting tasks on several widely used benchmark datasets. Our results show that in time series forecasting, Kernel-U-Net meets or outperforms baseline models, such as NLinear and PatchTST, in the majority cases.
In summary, the contributions of this work include:
  • We propose Kernel-U-Net, a general hierarchical and symmetrical framework that partitions input sequences into small slices and learns them at each layer of the network.
  • Kernel-U-Net generalizes the concept of kernels and provides convenience for composing particular models with custom kernels that follow the same design pattern. The parameter size is guaranteed in $O(log(L)^{2})$.
  • The computation complexity of Transformer module is $O(Llog(L))$ and can be reduced to $O(log(L)^{2})$ if it is only employed in the layers close to the latent vector.
  • Kernel-U-Net achieves results with MSE close to or better than the state-of-the-art methods on 6 (out of 7) time series data sets.
We conclude that Kernel-U-Net is a highly potential candidate for large-scale time series forecasting tasks. It is a general hierarchical and symmetrical framework with low complexity in terms of parameters. It reaches state-of-the-art results in the majority cases. With flexible customized kernels, it adapts well in high-paced development and test context.
2 Related works
Transformer
Transformer Vaswani et al. (2017) was initially introduced in the field of Natural Language Processing(NLP) on language translation tasks. It contains a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation. As it demonstrated outstanding performance on NLP tasks, many researchers follow this technique route. Bidirectional Encoder Representations from Transformers (BERT)Devlin et al. (2019) is designed to pre-train deep bidirectional representations from the unlabeled text by joint conditioning on both left and right context in all layers. Generative Pre-trained Transformer 3(GPT-3) pre-trained a massive transformer-based model using 175 billion parameters Brown et al. (2020) and achieved strong performance on different types of downstream natural language tasks without requiring any fine-tuning.
Later, Vision Transformers (ViTs) Dosovitskiy et al. (2021) ViT applies a pure transformer directly to sequences of image patches to classify the full image and outperformed CNN based method on ImageNetDeng et al. (2009). Swin Transformer Liu et al. (2021) proposed a hierarchical Transformer whose representation is computed with shifted windows. As a shifted window brings greater efficiency by limiting self-attention computation to non-overlapping local windows, it also allows cross-window connection. This hierarchical architecture has the flexibility to model at various scales and has linear computational complexity concerning image size.
In time series forecasting, the researchers were also attracted by transformer-based models. LogTrans Li et al. (2019) proposed convolutional self-attention by employing causal convolutions to produce queries and keys in the self-attention layer. It proposes also a LogSparse Transformer with only $O(L(logL)^{2})$ space complexity to break the memory bottleneck. Informer Zhou et al. (2021) has an encoder-decoder architecture that validates the Transformer-like model’s potential value to capture individual long-range dependency between long sequences. It proposes a ProbSparse self-attention mechanism to efficiently replace the canonical self-attention. It achieves the $O(LlogL)$ time complexity and $O(LlogL)$ memory usage on dependency alignments. PyraformerLiu et al. (2022a) simultaneously captures temporal dependencies of different ranges in a compact multi-resolution fashion. Theoretically, by choosing parameters appropriately, it achieves concurrently the maximum path length of $O(1)$ and the time and space complexity of $O(L)$ in forward pass.
Meanwhile, another family of transformer-based models combines the transformer with the traditional method in time series processing. Autoformer Wu et al. (2021) introduces an Auto-Correlation mechanism in place of self-attention, which discovers the sub-series similarity based on the series periodicity and aggregates similar sub-series from underlying periods. Frequency Enhanced Decomposed Transformer (FEDformer)Zhou et al. (2022) captures global properties of time series with seasonal-trend decomposition. The authors proposed Fourier-enhanced blocks and Wavelet-enhanced blocks in the Transformer structure to capture important time series patterns through frequency domain mapping.
Recently, YFormer Madhusudhanan et al. (2023) proposed a U-Net-like architecture that combines coarse and fine-grained features to capture long-range effects on all scale levels. PatchTST Nie et al. (2023) adopted the patches method from ViTs and proposed a time series transformer with a linear patch layer. It reaches state-of-the-art results on many time series forecasting benchmarks. The author claims Reduction in time and space complexity of learning transformer block from $O(L^{2})$ to $O((\frac{L}{S})^{2})$, where $L$ is sequence size and $S$ is stride size.
U-Net
U-Net is a neural network architecture designed primarily for medical image segmentation Ronneberger et al. (2015). U-net is composed of an encoder and a decoder. At the encoder phase, a long sequence is reshaped into small slices and reduced by kernels step by step into a latent vector. At the decoder stage, the vector is developed by kernels for generating an output with the same shape as the input. With the help of skip-connection between the encoder and decoder, U-Net can also capture and merge low-level and high-level information easily.
With such a neat structural design, U-Net has achieved great success in a variety of applications such as medical image segmentation Ronneberger et al. (2015), biomedical 3D-Image segmentation Çiçek et al. (2016), time series segmentation Perslev et al. (2019), image super-resolution Han et al. (2022) and image denoising Zhang et al. (2023). Its techniques evolve from basic 2D-U-Net to 3D U-Net Çiçek et al. (2016), 1D-U-Net Azar et al. (2022), Weighted Res-UNet Xiao et al. (2018) with residual structure and weighted attention mechanism, U-Net++ Zhou et al. (2018) with nested U-net inside each layer, UNet3+ Huang et al. (2020a) that allows full-scale skip connections incorporating low-level details with high-level semantics from feature maps in different scales and Swin-U-Net Cao et al. (2023) that only use transformer blocks at each layer of U-Net. Modification on U-Net layers has been studied by Selective Kernel U-Net Byra et al. (2020) which merges information from different sizes of convolutional kernels and Large Kernel U-Net Jia et al. (2022) which combines TransMorph layers for better performance in medical image segmentation.
In time series processing, U-time Perslev et al. (2019) is a U-Net composed of convolutional layers for time series segmentation tasks, and YFormer Madhusudhanan et al. (2023) is the first U-Net based model for time series forecasting task. In particular, YFormer applied transformer blocks on each layer of U-Net and capitalized on multi-resolution feature maps effectively.
Hierarchical and hybrid model
In time series processing, learning sequences of length $L$ requires $O(L)$ computations for recurrent models such as RNN, LSTM, GRU, and $O(L^{2})$ computations for transformer-based models. With ProbSparse Attention, the complexity can be reduced to $O(Llog(L))$. To balance the expressivity of complex modules and computational efficiency, researchers investigated hierarchical architecture and hybrid methods that merge different modules into the network.
For example, authors in Du et al. (2015) investigated a tree structure model made of bidirectional RNN layers and concatenation layers for skeleton-based action recognition, authors in Xiao and Liu (2016) stacked RNN and LSTM with Attention mechanism for semantic relation classification of text, authors in Kowsari et al. (2017) applied hierarchical LSTM and GRU was used for document classification. In time series processing, the authors in Hong and Yoon (2017) combine Deep Belief Network (DBN) and LSTM for sleep signal pattern classification, authors in LogTrans Li et al. (2019), Informer Zhou et al. (2021)) and Pyraformer Liu et al. (2022a) studied hierarchical or pyramidal transformers for time series forecasting.
Kernel-U-Net is a general hierarchical and symmetrical framework with low complexity in terms of parameter size. By replacing the convolutional kernel in classical U-Net, our model allows users to define custom kernels such as Attention, LSTM, or simple linear multiplication with a non-linear activation function for extracting features on complex time series. There are 3 advantages in Kernel-U-Net: 1) A small number of parameters: the total parameter of model is in order of $O(log(L)^{2})$, 2) Flexibility: the kernels can be replaced easily by arbitrary function that follows the same design mode so that non-linear kernels can be easily combined into it, 3) Computation efficiency: the computation complexity of transformer modules is reduce to $O(log(L)^{2})$ if they placed close to latent vector.
Algorithm 1 Kernel Wrapper
Input: kernel $f_{i}$, input segments $x$, input length $L_{in}$ and dimension $M_{in}$, output length $L_{out}$ , and dimension $M_{out}$
Output: Instance of Kernel Wrapper
1:  # Initiation of main kernel function $f$ is
2:  $f=$ $f_{i}(L_{in},M_{in},L_{out},M_{out})$
3:  # Holds variables for global operation
4:  _unet_skip_output = None
5:  _unet_skip_input = None
6:
7:  # Define the forward function as :
8:  Reshape $x$ into (-1, $L_{in}$,$M_{in}$)
9:  if _unet_skip_input is not None then
10:     $x=x+\_unet\_skip\_input$
11:  end if
12:  # Apply $f$ on $x$
13:  $y=f_{i}(x)$
14:  _unet_skip_output = $y$
15:  return $y$
3 Method
3.1 Problem Formulation
Let us note by $x\in R^{M\times N}$ the matrix which represents the multivariate time series dataset, where the first dimension $N$ represents the sampling time and the second dimension $M$ is the size of the feature. Let L be the length of memory or the look-back window, so $(x_{t+1,1},...,x_{t+L,M})$ (or for short, $(x_{t+1},...,x_{t+L}))$ is a slice of length $L$ of all features. It contains historical information about the system at instant $t$. We call $X\in R^{N\times M\times L}$ the trajectory matrix which represents the sliced historical data for each instant $t\in[0,N-L-1]$.
Time series forecasting.
In the case of time series forecasting, the dataset is composed of a time series of characteristics $x$ and future series $\hat{x}$. Let $x_{t}$ be the feature at the time step $t$ and $L$ the length of the look-back window. Given a historical data series $(x_{t+1},...,x_{t+L})$ of length $L$, time series forecasting task is to predict the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps. Then we can define the basic time series forecasting problem: $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})=f(x_{t+1},...,x_{t+L})$ , where $f$ is the function that predicts the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps based on a series $(x_{t+1},...,x_{t+L})$.
Algorithm 2 Kernel U-Net Encoder
Input: $L_{1},M_{1}$, $[,L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$, $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, $L_{h}$, $M_{h}$
Output: Instance of Kernel U-Net Encoder
1:  $L=[L_{1},L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$
2:  # Create first layer
[…]
4:  # Create intermediate layers
[…]
6:     layers.append(KernelWrapper($f_{i},L_{i},M_{i},1,M_{i+1}$))
7:  end for
8:  # Create last layer
9:  layers.append(KernelWrapper($f_{n+m},L_{n+m},M_{n+m},$ $L_{h},M_{h}$))
10:  encoder = nn.Sequential (*layers)
3.2 Kernel U-Net
Kernel U-Net follows the hierarchical structure of convolutional U-Net. It accepts a custom kernel to replace the linear kernel inside the original convolutional U-Net. More precisely, it repeatedly applies custom kernels on small slices of large-scale input data. In this context, Kernel U-Net avoids problems in processing long sequences such as gradient vanishing in recurrent models and quadradic complexity in transformers-based models. In the following paragraphs, we describe methods such as the hierarchical partition of the input sequence, and the creation of kernel wrapper, encoder, decoder, and U-net.
Hierarchical Partition of the input sequence.
Let us consider an input trajectory matrix $X_{t}\in R^{L\times M}$, where $L$ is lookback window size, $M$ is feature size and $t$ is the time step. Given a list of multiples $(L_{1},...,L_{n})$ for look-back window and $(M_{1},...,M_{m})$ for feature such that $L=\prod_{1}^{n}L_{i}$ and $M=\prod_{1}^{m}M_{i}$ We reshape $X_{t}$ as small slices $X_{t,i,j}\in R^{L_{1}\times M_{1}}$ such that the concatenation of all slices is equal to its original trajectory matrix:
$X_{t}=Concat([[X_{t,i,j}\forall i\in[1,\frac{L}{L_{1}}]]\forall j\in[1,\frac{M}{M_{1}}]])$.
The partitioned sequences are processed by kernels in the encoder. Since the operations in the decoder are symmetrical to those in the encoder, there will also be output $\hat{X}_{t+L}$ composed of generated segments $\hat{X}_{t+L,i,j}$.
Hierarchical processing with kernels in the encoder and decoder
Let us consider $X$ as a batch of trajectory matrix $X\in R^{B\times L\times M}$, where $B$ is the batch size. The hierarchical processing of Kernel U-Net consists of compressing an input sequence at the encoding stage and generating an output sequence at the decoding stage. By default, kernels reduce the dimension of input at each layer in an encoder and increase the dimension in a decoder.
At the stage of the encoder, the shape of input $X$ is $(B,L,M)$. Let us assume that $M_{hidden}$ is the unique dimension of the hidden vector at each intermediate layer to simplify the problem. We reshape $X$ into $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ then transpose it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ and reshape it into $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$ The kernel $f_{L_{1}}$ can now process the reshaped $X$ as a large batch of slices and output a vector $Z_{L_{1}}$ in shape of $(B*\prod_{2}^{m}*M_{i}*\prod_{2}^{n}L_{i},1,M_{hidden})$. After this operation, we reshape the output $Z_{L_{1}}$ into $(B*\prod_{2}^{m}M_{i}*\prod_{3}^{n}L_{i},L_{2},M_{hidden})$ for next layer. Iteratively, the encoder of Kernel U-Net processes all the multiples in the list $(M_{m},...,M_{2},L_{n},...,L_{2})$ in reverse order, and gives the final latent vector $Z$ in the shape of $(B,1,M_{hidden})$ as a batch of the latent vector.
At the stage of the decoder, the operations are reversed. We send a latent vector Z of shape $(B,1,M_{hidden})$ to decoder $f_{M_{m}}$ and get an output $Z_{M_{m}}$ in shape of $(B,M_{m},M_{hidden})$, then we reshape it into $(B*M_{m},1,M_{hidden})$ for the next kernels. At the end of this iteration over multiples in the list $(M_{m-1},...,M_{2},L_{n},...,L_{2},L_{1})$, we have finally a vector in the shape of $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$. The last operations are reshaping it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ , transposing it into the vector in shape $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ and reshaping it to shape $(B,L,M)$.
The algorithm is optimized in implementation because the for loop is expensive in the forward function of a PyTorch module. In our implementation, we nest the reshape and transpose operations in the kernel wrapper to avoid iteration over kernels. The complete algorithm is described in the next section.
Algorithm 3 Kernel U-Net Decoder
Input: $L_{h}$, $M_{h}$, $[L_{n},...,L_{2}]$, $[L_{n+m},...,L_{n+2}]$, $L_{1}$, $M_{1}$, $[f_{n+m},...,f_{n+2},$ $f_{n},...,f_{2}]$, $[M_{n+m},...,M_{n+2},$ $M_{n},...,M_{2}]$, $L_{1}$, $M_{1}$.
Output: Instance of Kernel U-Net Decoder
1:  # Create a list of multiples
[…]
3:  # Create first layer
4:  layers = [KernelWrapper($f_{n+m},$ $1,$ $M_{h},$ $L_{n+m},$ $M_{n+m}$)]
5:  # Create intermediate layers
6:  for  $i$ in $[n+m-1,...,n+2,n,...,2]$ do
[…]
8:  end for
9:  # Create last layer
10:  layers.append(KernelWrapper($f_{2},1,$ $M_{2},$ $L_{1},$ $M_{1})$)
11:  decoder = nn.Sequential (*layers)
Kernel Wrapper As described previously, Kernel U-Net accepts custom kernels that follow predefined design patterns. A kernel wrapper is an interface that executes the given kernel inside the encoder and decoder. It requires necessary parameters such as model kernel, input dimension, input length, output dimension, and output length and holds global variables such as skip connection vector. By default, a kernel reduces the length of a segment from $L_{i}$ to $1$ in the encoder and develops the length of a segment from $1$ to $L_{i}$ in the decoder. The kernel wrapper initiates an instance of a given kernel and calls it in the forward function.
Algorithm 4 Kernel U-Net
Input: $L_{1},M_{1}$, $[,L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$, $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, $L_{h}$, $M_{h}$
Output: Instance of Kernel U-Net
1:  # Initiation of encoder and decoder in K-U-N.
2:  encoder = KernelUNetEncoder(inputs)
3:  decoder = KernelUNetDecoder(inputs)
4:
5:  # Define the forward function as
6:  z = encoder(x)
7:  for  $i$ in $[1,...,n]$ do
8:     output = encoder.layers[i]._unet_skip_output
9:     decoder.layers[$n-i+1$]. _unet_skip_input = output
10:  end for
11:  y= decoder(z)
12:  return y
Creation of Kernel U-Net
Let us note input length $L$, input dimension $M$, Lists of multiples of look-back window $,[L_{2},$...,… , $L_{n}]$ and feature $,[L_{n+2},$...,… , $L_{n+m}]$ , size of segment-unit $L_{1}$ and feature-unit $M_{1}$ , a list of hidden dimension $[M_{2},...,M_{n},$ $M_{n+2},...,M_{n+m}]$, a list of kernels $[f_{1},f_{2},...,f_{n},$ $f_{n+2},...,f_{n+m}]$, latent vector length $L_{h}$ and dimension $M_{h}$.
Remark that the hidden dimension of the model is the size of intermediate output vectors within layers. It corresponds to channel size in a convolutional network. This dimension can be augmented for a larger passage of information. The kernel wrapper is described in algorithm 1. The creation of Kernel U-Net Encoder, Decoder, and complete U-Net are described in algorithms 2, 3, 4.
Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer Informer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.369 0.404 0.37 0.4 0.374 0.394 0.375 0.399 0.376 0.419 0.449 0.459 0.865 0.713
192 0.397 0.426 0.413 0.429 0.408 0.415 0.405 0.416 0.42 0.448 0.5 0.482 1.008 0.792
336 0.429 0.428 0.422 0.44 0.429 0.427 0.439 0.443 0.459 0.465 0.521 0.496 1.107 0.809
720 0.433 0.45 0.447 0.468 0.44 0.453 0.472 0.49 0.506 0.507 0.514 0.512 1.181 0.865
ETTh2
96 0.271 0.336 0.274 0.337 0.277 0.338 0.289 0.353 0.346 0.388 0.358 0.397 3.755 1.525
192 0.332 0.376 0.339 0.379 0.344 0.381 0.383 0.418 0.429 0.439 0.456 0.452 5.602 1.931
336 0.357 0.4 0.329 0.384 0.357 0.4 0.448 0.465 0.496 0.487 0.482 0.486 4.721 1.835
720 0.39 0.438 0.379 0.422 0.394 0.436 0.605 0.551 0.463 0.474 0.515 0.511 3.647 1.625
ETTm1
96 0.285 0.337 0.29 0.342 0.306 0.348 0.299 0.343 0.379 0.419 0.505 0.475 0.672 0.571
192 0.327 0.367 0.332 0.369 0.349 0.375 0.335 0.365 0.426 0.441 0.553 0.496 0.795 0.669
336 0.357 0.392 0.366 0.392 0.375 0.388 0.369 0.386 0.445 0.459 0.621 0.537 1.212 0.871
720 0.408 0.422 0.416 0.42 0.433 0.422 0.425 0.421 0.543 0.49 0.671 0.561 1.166 0.823
ETTm2
96 0.16 0.245 0.165 0.255 0.167 0.255 0.167 0.26 0.203 0.287 0.255 0.339 0.365 0.453
192 0.215 0.215 0.22 0.292 0.221 0.293 0.224 0.303 0.269 0.328 0.281 0.34 0.533 0.563
336 0.268 0.326 0.274 0.329 0.274 0.327 0.281 0.342 0.325 0.366 0.339 0.372 1.363 0.887
720 0.348 0.378 0.362 0.385 0.368 0.384 0.397 0.421 0.421 0.415 0.433 0.432 3.379 1.338
Electricity
96 0.129 0.225 0.129 0.222 0.141 0.237 0.14 0.237 0.193 0.308 0.201 0.317 0.274 0.368
192 0.148 0.244 0.147 0.24 0.154 0.248 0.153 0.249 0.201 0.315 0.222 0.334 0.296 0.386
336 0.168 0.263 0.163 0.259 0.171 0.265 0.169 0.267 0.214 0.329 0.231 0.338 0.3 0.394
720 0.205 0.299 0.197 0.29 0.21 0.297 0.203 0.301 0.246 0.355 0.254 0.361 0.373 0.439
Traffic
96 0.360 0.252 0.36 0.249 0.41 0.279 0.41 0.282 0.587 0.366 0.613 0.388 0.719 0.391
192 0.377 0.263 0.379 0.25 0.423 0.284 0.423 0.287 0.604 0.373 0.616 0.382 0.696 0.379
336 0.389 0.270 0.392 0.264 0.435 0.29 0.436 0.296 0.621 0.383 0.622 0.337 0.777 0.42
720 0.427 0.290 0.432 0.286 0.464 0.307 0.466 0.315 0.626 0.382 0.66 0.408 0.864 0.472
Weather
96 0.146 0.200 0.149 0.198 0.182 0.232 0.176 0.237 0.217 0.296 0.266 0.336 0.3 0.384
192 0.192 0.244 0.194 0.241 0.225 0.269 0.22 0.282 0.276 0.336 0.307 0.367 0.598 0.544
336 0.241 0.287 0.245 0.282 0.271 0.301 0.265 0.319 0.339 0.38 0.359 0.395 0.578 0.523
720 0.310 0.340 0.314 0.334 0.338 0.348 0.323 0.362 0.403 0.428 0.419 0.428 1.059 0.741
Table 1: Multivariate time series forecasting results with Kernel U-Net. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold and the second best results in underlined.
3.3 Custom Kernels
Linear kernel
The linear kernel is a simple matrix multiplication. Given $X$ of shape $(B,L_{i},M_{i})$, linear kernel $f_{i}$ reshape it into $(B,L_{i}*M_{i})$ and process it as follow:
$Z_{i}=f_{i}(X)=X*w+b$
, where $Z_{i}$ is output of kernel, $w\in R^{L_{i}M_{i}\times L_{i+1}M_{i+1}}$ is weight matrix and $b\in R^{L_{i+1}M_{i+1}}$ is bias vector. Remark that linear Kernel U-Net is an equivalent version to convolutional U-Net whose stride size is $L_{i}$. We also test a non-linear kernel that takes a non-linear activation function such as Tanh in the intermediate hidden layers.
Transformer kernel
The Vanilla Transformer is made of a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation.Wen et al. (2023)
Positional encoding (PE).
In a vanilla Transformer, the input sequence is added by a positional encoding vector given by:
$PE(t)=\left\{\begin{array}[]{ll}sin(\omega it)&ifi\%2=0\\ cos(\omega it)&ifi\%2=1\\ \end{array}\right.$
, where $\omega i$ is the frequency of waves generated at each dimension. By adding positional information to the input, the transformer can consider the same patterns that occur in different time steps as different patterns.
Multiple head attentions.
With the Query-Key-Value $(QKV)$ model, the scaled dot-product attention used by Transformer is given by
$Attention(Q,K,V)=Softmax(\frac{QK^{T}}{\sqrt{D_{k}}}))V$
, where queries $Q\in R^{N\times D_{k}}$ , keys $K\in R^{M\times D_{k}}$ , values $V\in R^{M\times D_{v}}$ , $N,M$ denote the lengths of queries and keys (or values), and $D_{k},D_{v}$ denote the dimensions of keys (or queries) and values. Remark that in the context of self-attention, $Q=K=V=inputdata$.
The multiple attention block contains repeated heads and can be given by
$MultiHeadAttn(Q,K,V)=Concat(h_{1},...,h_{H})W^{O}$
, where $h_{i}=Attention(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}).$
LSTM kernel
The Long Short-term memory contains 3 gates: Forget Gate $f_{t}$, Input Gate $i_{t}$, Output Gate $o_{t}$. Each gate has a weight matrix $W$ and a bias vector $b$ for the current input $x_{t}$ and the previous hidden state $h_{t-1}$. The next cell state $C_{t}$ is updated by the current activated cell state $C^{\prime}_{t}$. The next hidden state $h_{t}$ is updated by the output gate and the next activated cell state. The formulation of LSTM is as follows:
$f_{t}=\sigma(W_{f}[h_{t-1},x_{t}]+b_{f})$
$i_{t}=\sigma(W_{i}[h_{t-1},x_{t}]+b_{i})$
$C^{\prime}_{t}=tanh(W_{C}[h_{t-1},x_{t}]+b_{C})$
$C_{t}=f_{t}C_{t-1}+i_{t}C^{\prime}_{t}$
$o_{t}=\sigma(W_{o}[h_{t-1},x_{t}]+b_{o})$
$h_{t}=o_{t}tanh(C_{t})$
Methods K-U-Net PatchTST Nlinear Dlinear FEDformer Autoformer    LogTrans
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.052 0.176 0.055 0.179 0.053 0.177 0.056 0.18 0.079 0.215 0.071 0.206 0.283 0.468
192 0.067 0.202 0.071 0.215 0.069 0.204 0.071 0.204 0.104 0.245 0.114 0.262 0.234 0.409
336 0.074 0.22 0.076 0.22 0.081 0.226 0.098 0.244 0.119 0.27 0.107 0.258 0.386 0.546
720 0.074 0.223 0.087 0.232 0.08 0.226 0.189 0.359 0.127 0.28 0.126 0.283 0.475 0.629
ETTh2
96 0.114 0.262 0.129 0.282 0.129 0.278 0.131 0.279 0.128 0.271 0.153 0.306 0.217 0.379
192 0.149 0.304 0.168 0.328 0.169 0.324 0.176 0.329 0.185 0.33 0.204 0.351 0.281 0.429
336 0.158 0.319 0.171 0.336 0.194 0.355 0.209 0.367 0.231 0.378 0.246 0.389 0.293 0.437
720 0.178 0.34 0.223 0.38 0.225 0.381 0.276 0.426 0.278 0.42 0.268 0.409 0.218 0.387
ETTm1
96 0.026 0.122 0.026 0.121 0.026 0.122 0.028 0.123 0.033 0.14 0.056 0.183 0.049 0.171
192 0.039 0.15 0.039 0.15 0.039 0.149 0.045 0.156 0.058 0.186 0.081 0.216 0.157 0.317
336 0.051 0.173 0.053 0.173 0.052 0.172 0.061 0.182 0.071 0.209 0.076 0.218 0.289 0.459
720 0.068 0.202 0.073 0.206 0.073 0.207 0.08 0.21 0.102 0.25 0.11 0.267 0.43 0.579
ETTm2
96 0.061 0.181 0.065 0.187 0.063 0.182 0.063 0.183 0.063 0.189 0.065 0.189 0.075 0.208
192 0.085 0.22 0.093 0.231 0.09 0.223 0.092 0.227 0.102 0.245 0.118 0.256 0.129 0.275
336 0.104 0.25 0.121 0.266 0.117 0.259 0.119 0.261 0.13 0.279 0.154 0.305 0.154 0.302
720 0.149 0.303 0.171 0.322 0.17 0.318 0.175 0.32 0.178 0.325 0.182 0.335 0.16 0.321
Table 2: Univariate time series forecasting results with Kernel U-Net. The prediction lengths $T$ are in {96, 192, 336, 720} for all datasets. We note the best results in bold and the second best results in underlined.
The LSTM Kernel used in our experiment contains an LSTM cell and a linear layer that takes all hidden states of a sequence as input and gives a vector as output. The experimental result is shown in the section 4
4 Experiments and Results
Datasets.
We conducted experiments on our proposed Kernel U-Net on 7 public datasets, including 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Weather, Traffic and Electricity. These datasets have been benchmarked and publicly available onWu et al. (2021) and the description of the dataset is available in Zeng et al. (2023)).
Here, we followed the experiment setting in Zeng et al. (2023) and partitioned the data into 12, 4, and 4 months for training, validation, and testing respectively for the ETT dataset. The data is split into [0.7,0.1,0.2] for training, validation, and testing for dataset Weather, Traffic, and Electricity.
Baselines and Experimental Settings.
In the time series forecasting task, we follow the experiment setting in NLinear Zeng et al. (2023), which take $L={336,720}$ step historical data as input and forecast $T\in{96,192,336,720}$ step value in the future. We replace the last value normalization with mean value normalization for univariate forecasting in the ETT dataset. We include five recent methods: PatchTSTNie et al. (2023), NLinear, DLinearZeng et al. (2023), FEDformer Zhou et al. (2022), Autoformer Wu et al. (2021), Informer Zhou et al. (2021)and LogTrans Li et al. (2019). We merge the result reported in Nie et al. (2023) and Zeng et al. (2023) for taking their best in a supervised setting.
Following previous works Wu et al. (2021), we use Mean Squared Error (MSE) and Mean Absolute Error (MAE) as the core metrics to compare performance for Forecasting problems. We take a look back at Windows 336 and 720 in experiments. The list of multiples is respectively [3, 4, 4, 5] and [4, 4, 7] and the segment-unit input length is 3. The input dimension is 1 as we follow the channel-independent setting in Zeng et al. (2023). The hidden dimensions are 128 or 256 for multivariate tasks and 5 or 30 for univariate tasks depending on of dataset. We use the Random Erase Zhong et al. (2020) from image processing for data augmentation. The learning rate is selected in [0.00005, 0.001] depending on the dataset. The training epoch is 300 and the patience is 30 in general.
Instance Normalization and Mean Normalization Instance normalization Ulyanov et al. (2017) Kim et al. (2021) contains a step of substruction with the average value of the input sequence and a division of its standard deviation. It computes reversely the multiplication of its standard deviation and the addition of the average value after the output of the model. Mean normalization removes the division and multiplication step of the standard deviation of instance normalization to avoid introducing noises in the small dataset. We followed the mean normalization as the preprocessing step and compared it with instance normalization for selected models.
Model Variants.
We propose 4 kernels and 5 versions of kernel U-Net for experiments. The kernels tested are Linear, Linear with 1 hidden layer and non-linear activation, Transformer, and LSTM.
  • Linear U-Net is a baseline method that contains a list of linear kernels.
  • Linear-1-Hidden U-Net replaces 2 kernels close to latent vector to kernels with hidden layer and nonlinear activation.
  • Linear-5-Hidden U-Net replaces all linear kernels with non-linear kernels.
  • Transformer U-Net replaces 2 kernels close to latent vector to transformer kernels and keeps the rest as linear kernels.
  • LSTM U-Net replaces 2 kernels close to latent vector to LSTM kernels and keeps the rest as linear kernels.
The result shows that the performance of Kernel U-Net reaches or outperforms the state-of-art methods in univariate setting (Table 2) and reaches similar or better results in multivariate setting (Table 1).
5 Conclusion
In this paper, we propose Kernel U-Net, a symmetric and hierarchical framework that can effectively process long temporal dependencies with $O(log(L)^{2})$ number of parameters. Specifically, Kernel U-Net cut input sequences into small slices at each layer in the network for applying customized kernels.
In the future, we hope to develop more kernels that adapt well to particular datasets. We also hope that Kernel U-Net can be useful for other time series tasks such as classification or anomaly detection.